# Building Trustworthy AI by Addressing its 16+2 Desiderata with Goal-Directed Commonsense Reasoning

### v1

@@ Abstract
Current advances in AI and its applicability have highlighted the need to ensure its trustworthiness for legal, ethical, and even commercial reasons. Sub-symbolic machine learning algorithms, such as the LLMs, simulate reasoning but hallucinate and their decisions cannot be explained or audited (crucial aspects for trustworthiness). On the other hand, rule-based reasoners, such as Cyc, are able to provide the chain of reasoning steps but are complex and use a large number of reasoners. We propose a middle ground using s(CASP), a goal-directed constraint-based answer set programming reasoner that employs a small number of mechanisms to emulate reliable and explainable human-style commonsense reasoning. In this paper, we explain how s(CASP) supports the 16 desiderata for trustworthy AI introduced by Doug Lenat and Gary Marcus (2023), and two additional ones: inconsistency detection and the assumption of alternative worlds. To illustrate the feasibility and synergies of s(CASP), we present a range of diverse applications, including a conversational chatbot and a virtually embodied reasoner.
@@ 1 Introduction
Although incredible advancements in artificial intelligence (AI) have been made, robustness in AI systems is still lacking. The recent AAAI study by ? (?) found that 76% of AI professionals surveyed believed that current approaches were “unlikely” or “very unlikely” to yield true artificial general intelligence (AGI). Despite that, the predominant approach towards AI remains pouring more resources—data, training time, and GPUs—into generative AI in order to produce marginally better results. We argue that what is necessary is not larger and more robust neural networks, but rather the encoding of commonsense knowledge alongside a capable reasoner. In this paradigm, deep learning becomes a tool for a reasoner to interface with us, not an entity in and of itself.
@@
Researchers have agreed since the inception of AI (?) that commonsense knowledge and reasoning are important for the field. There are two fundamental components in commonsense reasoning: the knowledge base, containing all the commonsense knowledge about the problem domain, and the commonsense reasoner, which processes that knowledge to draw new conclusions.
@@
Sub-symbolic vs. symbolic AI: The first part, the knowledge base, almost universally poses a problem for modern deep learning methods. In deep learning, knowledge is extracted from huge amounts of data. It is impossible to perform any reasonable quality assurance at the required scale . This problem is being tackled by improving the ability to train on small datasets, see survey by ? (?), and through the use of services like Amazon’s Mechanical Turk to label large quantities of data with inexpensive labor. Using data that is not labeled or evaluated for quality tends to lead to difficulty with reasoning in these systems. In some domains, a strong foundational knowledge base is almost inherently machine-usable, such as in physics-informed machine learning, see survey by ? (?), where commonsense knowledge takes the form of math equations. However, in others, such as the processing of natural language, the translation from knowledge to reasoning is more difficult (?). Deep learning systems are incredibly robust to ingesting large quantities of data, but they do not reason and are prone to errors that would not be made by a system with sound commonsense knowledge. On the other hand, logic-based symbolic reasoners struggle with representing large amounts of knowledge that can be processed efficiently. A combination of deep learning and logic-based reasoning—a neuro-symbolic system—perform better on key metrics like explainability without sacrificing performance, see survey by ? (?). Overall, systems that combine a robust knowledge base with a reasoner designed to utilize it perform the best, though the problem is far from solved.
@@
Cyc, a machine reasoner: One of the oldest attempts to use commonsense knowledge is the Cyc project. Cyc seeks to encode all commonsense knowledge into a knowledge base following its own proprietary ontology (?). For four decades the Cyc knowledge base has grown, gaining hand-crafted and hand-categorized axioms of knowledge covering every imaginable subject and ambiguity in the hopes of using that knowledge to cultivate true human-like intelligence. Cyc contains not just facts (like that the sky is blue) but also relations that are self-explanatory for humans, such as that if a person is a citizen of the same country as another, they have the same heads of state but different heads on their bodies. Because the Cyc knowledge base is proprietary, it is difficult to say how extensive the knowledge base is, but it is safe to say that 40 years has made it vast, requiring a large amount of reasoners on top of it to extract useful knowledge. These 1,100+ reasoners were designed with Cyc in mind and are adept at navigating it to solve different problems.
@@
However, outside of Cycorp the reception of Cyc has been lukewarm. Cyc’s philosophy has not pervaded AI research, possibly due to Cyc being a closed system, and deep learning based AI research has moved to the forefront. Many of Cyc’s critics use it as an example to prove that symbolic AGI is doomed for failure. However, as we look to the next steps for AGI beyond deep learning, other symbolic approaches solve some of its greatest weaknesses and show promise for the future. We posit that purely deep learning systems require symbolic intervention to improve their trustworthiness. Explainability is fundamentally important for AI systems that are trustworthy enough for critical use (?).
@@
LLMs and Trustworthiness: A recent paper written by ? (?) discusses what is needed for truly trustworthy AI. The paper argues that Large Language Models (LLMs) are untrustworthy, unstable, and brittle, something that has continued to prove true as LLMs have become more and more ubiquitous. ? (?) posit that generative AI systems fail because they do not reason, but rather are very sophisticated pattern matching algorithms. They argue that LLMs would need to pair up with a system like Cyc to become trustworthy. However, Cyc is a proprietary and complex system.
@@
Answer Set Programming: For symbolic reasoning systems to succeed, simplicity must be paramount. The answer set programming (ASP) paradigm has shown how commonsense knowledge can be represented and reasoned over in an elegant way (???). Implementations of ASP such as CLINGO ? (?), however, have a problem, namely, that they are based on grounding the program and then using a SAT solver to find answer sets. Thus, they cannot scale and are primarily used to solve combinatorial problems rather than large knowledge-based reasoning problems. To alleviate this, we propose to model commonsense reasoning with goal-directed ASP engines, which do not ground programs and support constructive negation, giving rise to what we have termed goal-directed commonsense reasoning. In particular, we propose the use of s(CASP), of which we provide a brief review in section 2.
@@
Addressing 16+2 Desiderata: This paper aims to demonstrate how s(CASP) incorporates the 16 desiderata established by ? (?) as capabilities required by general AI to be trustworthy. We also discuss additional desiderata that a (goal-directed) commonsense reasoning system must satisfy: (i) being able to specify inconsistencies via global constraints, and (ii) being able to incorporate assumptions through multiple possible worlds. The s(CASP) system incorporates these additional desiderata as well. Given the simplicity of s(CASP), the fact that it is open-sourced, and that it supports all the desiderata, we present it as a publicly available system for building AGI. These 16+2 desiderata are reviewed in Section 3 with a brief description, related s(CASP) capabilities, and a discussion of related work. The main contribution of this paper is to organize the capabilities of s(CASP), and present some of its initial applications, in terms of the desiderata for trustworthy AI to build solutions to the problems it faces.
@@ 2 Background: s(CASP)
s(CASP), by ? (?), is a novel non-monotonic reasoner that evaluates Constraint Answer Set Programs without a grounding phase either before or during execution. s(CASP) supports predicates and thus retains logical variables (and constraints) both during the execution and in the answer sets. The operational semantics of s(CASP) relies on backward chaining, which is intuitive to follow and lends itself to generating explanations that can be translated into natural language (NL). The execution of an s(CASP) program is goal-directed (i.e., starts with a query), and returns partial stable models: the subsets of the stable models, defined by ? (?), which include only the (negated) literals necessary to support the initial query. s(CASP) has been already applied in relevant works related to the representation of commonsense reasoning:
VECSR: Virtually Embodied Common Sense Reasoning (VECSR) by ? (?) is a system that accepts high-level tasks and uses s(CASP) to plan and execute those tasks in a context-aware fashion in VirtualHome, a fully-embodied simulated environment.
Reliable Chatbots: In the Amazon Alexa Socialbot Grand Challenge 4, ? (?) developed a conversational AI chatbot based on two natural language understanding systems by ? (?). Going further, ? (?) used LLMs to translate natural language into predicates (and vice versa) and employ commonsense reasoning based on s(CASP) to build a reliable task-bot as well as a socialbot to hold a social conversation with a human.
Event Calculus (EC): A reasoner on EC, by ? (?), has been used by ? (?) to model real-world avionics systems, verify (timed) properties, and identify gaps in system requirements, and by ? (?) to model medical devices.
Justification in NL: Has been used by ? (?) and ? (?) to bring Explainable Artificial Intelligence (XAI) principles to expert knowledge systems.
ILP systems: Inductive Logic Programming (ILP) systems generate ASP programs from data. This includes LIME-FOLD proposed by ? (?) and its extensions FOLD-RM by ? (?), FOLD-R++ by ? (?) and FOLD-SE by ? (?).
s(LAW): An administrative and judicial discretion reasoner by ? (?), which allows models legal rules involving ambiguity to infer conclusions and provides natural language justifications for those conclusions.
Spatial reasoner: Implemented by ? (?), it models dynamic information and restrictions in Building Information Modeling (BIM) and paves the way to using logic-based methodologies such as model refinement.
@@ 3 Addressing 16+2 desiderata with s(CASP)
The 16 desiderata for trustworthy AI are described in detail by ? (?), and so this section will only briefly summarize them. For each of the 16+2 desiderata we provide examples of how s(CASP) systems perform against them. To accomplish some desiderata we describe how s(CASP) successfully combines many of the benefits of both symbolic and neural AI systems. Additionally, we discuss the accomplishments of state-of-the-art systems, primarily Cyc and deep learning systems.
[…]
@@
Explainability is a commonly desired trait of AI, stating that any trustworthy AI must be able to explain any solution found. Its importance is reflected in the growing field of XAI (?).
s(CASP)
@@
Goal-directed execution means that one poses queries that are answered by an SLD-resolution-like process, similar to Prolog, and therefore, a proof is explicitly constructed. This proof constitutes not only the explanation but also the justification of a given decision (positive or negative). In s(CASP), the justification for any query is the trace of the successful derivation, or in cases where the query fails, is the trace of the negated query. Moreover, the framework developed by ? (?) allows the generation of justifications in natural language and navigable HTML files.
Thus, explainability is native to s(CASP) and has been exploited in applications such as the physician advisor for chronic heart failure management by ? (?).
@@
Discussion: One of the great weaknesses of deep learning systems is that they are unexplainable black boxes, where it is difficult to justify a solution. The closest LLMs get is that it can be asked “why” an answer was given, but the explanation provided may be for a different alternative solution or the justification itself may be a hallucination.
[…]
@@
A foundational aspect of trustworthy AI is its capacity for human-like deduction, encompassing logical reasoning (e.g., modus ponens), arithmetic operations, exhaustive search, and the identification of contradictions or redundancies. This includes understanding logical connectives, in particular recognizing different forms of negation, i.e., not being able to conclude p is different from being able to conclude p is false.
s(CASP)
[…]
@@
Discussion: Prior research has found that these deductive (along with inductive and abductive) reasoning skills are somewhat lacking in LLMs (??). Cyc, on the other hand, provides mixed results. Cyc has over 1,000 reasoners for deductions that are each optimized for different kinds of knowledge, contexts, and solution finding. This provides a wide breadth of deduction ability, however there are disadvantages associated with having such a large number of reasoning engines. Namely, processing power required by running multiple reasoners and overhead from meta-reasoners deciding which reasoners are needed. There is a trade-off between accuracy and speed associated with symbolic or neural methods.
[…]
Effective inductive reasoning is essential for navigating complex, uncertain environments. It involves generalizing from specific observations, such as inferring species traits, and making temporal projections where the probability of a fact holding over time often follows predictable decay curves (e.g., linear, Gaussian). Such reasoning supports adaptive decision-making, despite the inherent risk of error. This type of reasoning is the foundation for explanation-based learning, which seeks to learn more with less data based on inducing general rules from specific examples (?).
s(CASP)
Inductive Logic Programming (ILP), see survey by ? (?), is a sub-field of machine learning that learns interpretable logic programs from labeled data and background knowledge. Unlike statistical models, ILP uses symbolic representations and logical inference to derive hypotheses that generalize observations in a human-understandable way. This makes ILP valuable in domains requiring transparency, prior knowledge integration, and formal reasoning, such as bio-informatics, legal systems, and explainable AI.
The FOLD family of algorithms extends ILP by enabling the learning of default rules with exceptions, capturing common patterns of non-monotonic reasoning. Given the following data, the initially learned rule is flies(X) :- bird(X) due to the data on “tweety”. Later, from the data on “pengu” learns that penguins are an exception to the rule, so the rule flies(X) is modified, by adding not ab(X).
@@
bird(tweety). flies(tweety).
bird(pengu). penguin(pengu). -flies(pengu)
@@
flies(X) :- bird(X), not ab(X).
ab(X) :- penguin(X).
They have been used to encode various real-world data into s(CASP) programs. All versions of the FOLD algorithms, which we already enumerated in Section 2, inductively learn rule-sets from small datasets, provide a comparable accuracy to deep learning systems, and are completely explainable, thanks to the s(CASP) justification framework.
@@
Analogy is the ability to map similarity between objects of often disparate meaning. This is defined as a representational mapping from a known “source” domain to a novel “target” domain (?). In natural language, this commonly comes in the form of analogical phrases (“life is like a box of chocolates”). However, analogical reasoning is also used for various logic problems (such as Raven’s Progressive Matrices, or RPMs). Analogical reasoning of both kinds have been historically difficult for AI systems.
s(CASP)
@@
Symbolic systems have been used to solve the analogical reasoning problem with various degrees of success (?). Because these other successful systems (??) are often predicated on turning analogical problems into predicates that can then be reasoned over, s(CASP) could be used as a more modern reasoner with at least as much success.
@@
In the social chatbot application by ? (?), analogical reasoning is achieved by an integration of s(CASP) and LLMs. If the LLM parser recognized that the current topic is about the movie Titanic, an analogical reasoning will be triggered to find, for example, Catch Me If You Can as a candidate for the next topic, since Leonardo DiCaprio leads in both.
relevant_topic(A, B, has_same_person’) :-
movie(A), movie(B), person(C),
act_in(C, A), act_in(C, B).
Additionally, LLMs also help generate analogical reasoning rules or predicates:
@@
relevant_topic(’Batman Dark Knight Rises’,
Titanic’, hero sacrifices himself for others’).
Thus, s(CASP) can “solve” both analogical riddles and analogical phrases.
@@
Discussion: LLMs have been found to perform well at responding to natural language analogical phrases (?). However, in analogical reasoning, LLMs often make mistakes different than what humans would make due to a lack of deeper understanding. The ability to replicate responses to analogical queries is valuable but insufficient for true trustworthy AI. Alternatively, symbolic approaches can achieve good results on logical puzzles like RPMs, but lack the ability to integrate with natural language or make logical leaps like analogical phrases require without careful data population. In that area, Cyc likely does well on analogical phrases contained in its database.
[…]

### v2

@@ Abstract
Current advances in AI and its applicability have highlighted the need to ensure its trustworthiness for legal, ethical, and even commercial reasons. Sub-symbolic machine learning algorithms, such as the LLMs, simulate reasoning but hallucinate and their decisions cannot be explained or audited (crucial aspects for trustworthiness). On the other hand, rule-based reasoners, such as Cyc, are able to provide the chain of reasoning steps but are complex and use a large number of reasoners. We propose a middle ground using s(CASP), a goal-directed constraint-based answer set programming reasoner that employs a small number of mechanisms to emulate reliable and explainable human-style commonsense reasoning. In this paper, we explain how s(CASP) supports the 16 desiderata for trustworthy AI introduced by Doug Lenat and Gary Marcus (2023), and two additional ones: inconsistency detection and the assumption of alternative worlds. To illustrate the feasibility and synergies of s(CASP), we present a range of diverse applications.
keywords = , breaklines = true, breakatwhitespace=true, alsoletter=’,., basewidth = 0.42em, upquote = true, basicstyle = , moredelim = *[s][](), moredelim = *[s][]:-., literate = :-:–2 ?-?–2 .>.#>4 .<.#<4 .=.#=4 $\neq$4 .>=.#>=5 .=<.#=<5 ==2 moredelim = *[s][]:-., basicstyle = , moredelim = *[s][](), numbers=none, xleftmargin=0cm, basewidth = 0.42em, xleftmargin=0.5cm, numberstyle=, numbers=left, stepnumber=1, mathescape = true, keywords = , upquote = true, basicstyle = , basewidth = 0.42em, moredelim = *[s][]:-., moredelim = [s][](), commentstyle = , morecomment=[l]%, literate = .>.#>4 .<.#<4 .=.#=4 $\neq$4 .>=.#>=5 .=<.#=<5 —1
alexisrenee1@gmail.com yankai.zeng@utdallas.edu huaduo.wang@utdallas.edu joaquin.arias@urjc.es gupta@utdallas.edu
1University of Texas at Dallas, Texas, USA
2CETINIA, Universidad Rey Juan Carlos, Madrid, Spain
@@ 1 Introduction
Although incredible advancements in artificial intelligence (AI) have been made, robustness in AI systems is still lacking. The recent AAAI study by Rossi & et al. (2025) found that a “majority of respondents (76%) assert that ‘scaling up current AI approaches’ to yield AGI is ‘unlikely’ or ‘very unlikely’ to succeed, suggesting doubts about whether current machine learning paradigms are sufficient for achieving general intelligence”. Despite that, the predominant approach towards AI remains pouring more resources, such as data, training time, and GPUs, into generative AI in order to produce marginally better results. We argue that what is necessary is not larger and more robust neural networks, but rather the encoding of commonsense knowledge alongside a capable reasoner. In this paradigm, deep learning becomes a tool for a reasoner to interface with, not an entity in itself.
@@
Researchers have agreed since the inception of AI that commonsense knowledge and reasoning are important for the field. There are two fundamental components in commonsense reasoning: the knowledge base, containing all the commonsense knowledge about the problem domain, and the commonsense reasoner, which processes that knowledge to draw new conclusions.
@@
Sub-symbolic vs. symbolic AI: The first part, the knowledge base, almost universally poses a problem for modern deep learning methods. In deep learning, knowledge is extracted from huge amounts of data. It is impossible to perform any reasonable quality assurance at the required scale. This problem is being tackled by improving the ability to train on small datasets and through the use of services like Amazon’s Mechanical Turk to label large quantities of data with inexpensive labor. Using data that is not labeled or evaluated for quality tends to lead to difficulty with reasoning in these systems. In some domains, a strong foundational knowledge base is almost inherently machine-usable; however, in others the translation from knowledge to reasoning is more difficult. Deep learning systems are incredibly robust to ingesting large quantities of data, but they do not reason and are prone to errors that would not be made by a system with sound commonsense knowledge. On the other hand, logic-based symbolic reasoners struggle with representing and efficiently processing large amounts of knowledge. A combination of deep learning and logic-based reasoning, a neuro-symbolic system, performs better on key metrics like explainability without sacrificing performance, see survey by Yu et al. (2023). Overall, systems that combine a robust knowledge base with a reasoner designed to utilize it, perform the best. The problem is far from solved though, as brittleness of the logic-based knowledge base and commonsense reasonng remains a challenge.
@@
Cyc, a machine reasoner: One of the oldest attempts to use commonsense knowledge is the Cyc project. Cyc seeks to encode all commonsense knowledge into a knowledge base following its own proprietary ontology (Lenat, 1995). For four decades the Cyc knowledge base has grown, gaining hand-crafted and hand-categorized axioms of knowledge covering every imaginable subject and ambiguity in the hopes of using that knowledge to cultivate true human-like intelligence.
@@
However, outside of Cycorp the reception of Cyc has been lukewarm. Cyc’s philosophy has not pervaded AI research, possibly due to Cyc being a closed system, and deep learning-based AI research has moved to the forefront. Many of Cyc’s critics use it as an example to prove that symbolic AI is doomed for failure. However, as we look to the next steps for AI beyond deep learning, other symbolic approaches solve some of its greatest weaknesses and show promise for the future. We posit that purely deep learning systems require symbolic intervention to improve their trustworthiness. Explainability is fundamentally important for AI systems that are trustworthy enough for critical use.
@@
LLMs and Trustworthiness: Recent work by Lenat & Marcus (2023) discusses what is needed for truly trustworthy AI. The paper argues that Large Language Models (LLMs) are untrustworthy, unstable, and brittle, something that has continued to prove true as LLMs have become more and more ubiquitous. They posit that generative AI systems fail because they do not reason, but rather are very sophisticated pattern matching algorithms. They argue that LLMs would need to pair up with a system like Cyc to become trustworthy. However, Cyc is a complex and proprietary system.
@@
Answer Set Programming: For symbolic reasoning systems to succeed, simplicity must be paramount. The answer set programming (ASP) paradigm has shown how commonsense knowledge can be represented and reasoned over in an elegant way (Brewka et al., 2011; Gelfond & Kahl, 2014; Gupta et al., 2022). Implementations of ASP such as the CLINGO system (Gebser et al., 2014), however, have a problem, namely, that they are based on grounding the program and then using a SAT solver or a procedure inspired by SAT to find answer sets. Thus, they do not scale well and are primarily used to solve combinatorial problems rather than large knowledge-based reasoning problems. To alleviate this, we propose to model commonsense reasoning with goal-directed ASP engines, which do not ground programs and support constructive negation, giving rise to what we have termed goal-directed commonsense reasoning. In particular, we propose the use of s(CASP), of which we provide a brief review in Section 2.
@@
Addressing 16+2 Desiderata: This paper provides a structured argument that s(CASP) plausibly satisfies the desiderata established by Lenat & Marcus (2023) as capabilities required by general AI to be trustworthy. We also discuss additional desiderata that a (goal-directed) commonsense reasoning system must satisfy: (i) being able to specify inconsistencies via global constraints, and (ii) being able to incorporate assumptions through multiple possible worlds. The s(CASP) system incorporates these additional desiderata. Given the simplicity of s(CASP), the fact that it is open-sourced, and that it can support all the desiderata, we present it as a publicly available system for building trustworthy AI in conjunction with machine learning. These 16+2 desiderata are reviewed in Section 3 with a brief description, related s(CASP) capabilities, and a discussion of related work. The main contribution of this paper is to organize the capabilities of s(CASP), and present some of its applications in terms of the desiderata for trustworthy AI.
@@ 2 Background: s(CASP)
s(CASP), by Arias et al. (2018), is a novel non-monotonic reasoner that evaluates Constraint Answer Set Programs without a grounding phase either before or during execution. s(CASP) supports predicates and thus retains logical variables (and constraints) both during the execution and in the answer sets. The operational semantics of s(CASP) rely on backward chaining, which is intuitive to follow and lends itself to generating explanations that can be translated into natural language (NL). The execution of an s(CASP) program is goal-directed (i.e., starts with a query), and returns partial stable models: the subsets of the stable models, defined by Gelfond & Lifschitz (1988), which include only the positive or negative literals necessary to support the initial query. Any computed bindings for unbound variables in the query are also returned. The s(CASP) system supports first order predicates and terms, constructive negation, and coinduction. More details on s(CASP) are not included due to lack of space, but they can be found elsewhere (Arias et al., 2018; Gupta et al., 2022). s(CASP) has been used for many applications related to the representation of commonsense reasoning (including commercial ones as in Forsante Oy (2024)). We outline the ones below that are crucial for this paper:
VECSR: Virtually Embodied Common Sense Reasoning (VECSR) by Tudor et al. (2025) is a system that accepts high-level tasks and uses s(CASP) to plan and execute those tasks in a context-aware fashion in VirtualHome, a fully-embodied simulated environment.
Reliable Chatbots: In the Amazon Alexa Socialbot Grand Challenge 4, Basu et al. (2021b) developed a conversational AI chatbot based on two natural language understanding systems by Basu et al. (2021a). Going further, Zeng et al. (2024) used LLMs to translate natural language into predicates (and vice versa) and employ commonsense reasoning based on s(CASP) to build a reliable task-bot as well as a socialbot to hold a social conversation with a human.
Event Calculus (EC): A reasoner on EC by Arias et al. (2022a) has been used by Hall et al. (2021) to model real-world avionics systems, verify timed properties, and identify gaps in system requirements, and by Vašíček et al. (2024) to model a medical device.
Justification in NL: Has been used by Arias et al. (2020) and Chen et al. (2016) to bring eXplainable Artificial Intelligence (XAI) principles to expert knowledge systems.
ILP systems: Inductive Logic Programming (ILP) systems generate logic programs from data. This includes the FOLD family of algorithms, the latest of which is FOLD-SE by Wang & Gupta (2024) that learns stratified answer set programs.
s(LAW): An administrative and judicial discretion reasoner by Arias et al. (2024a), which allows modeling of legal rules involving ambiguity to infer conclusions and provides natural language justifications for those conclusions.
Spatial reasoner: Implemented by Arias et al. (2022c), it models dynamic information and restrictions in Building Information Modeling (BIM) and paves the way to using logic-based methodologies such as model refinement.
@@ 3 Addressing 16+2 desiderata with s(CASP)
Table 1: Desiderata support by Cyc, Deep-Learning, and s(CASP). Including s(CASP) features with citation.
Cyc DL s(CASP)
Feature / Tool
1 Explanation Justification Trees[1]
1 Deduction ! Base s(CASP)[2]
1 Induction ! ! Default Rules + FOLD[3]
1 Analogy ! ! ! Proposed LLM + s(CASP)
1 Abductive Reasoning Abducibles[2]
1 Theory of Mind Chatbot[4]
1 Quantifier-fluency Goal-Directed Execution[2]
1 Modal-fluency Deontic Logic[5]
1 Defeasibility Default Rules + FOLD[3]
1 Pro and Con Arguments ASP Multiple Worlds[6]
1 Contexts VECSR[7]
1 Meta-knowledge and Meta-reasoning Justification Trees[1]
1 Explicitly Ethical ! Constraints[2]
1 Sufficient Speed ! VECSR[7]
1 Sufficiently Lingual and Embodied Chatbot[4]
1 Broadly and Deeply Knowledgeable ! FOLD, Chatbot, VECSR[3,4,7]
1 Inconsistency Detection ! Constraints[2]
1 Multiple Possible Worlds ! ASP Multiple Worlds[6]
[1] Arias et al. (2020), [2] Arias et al. (2018), [3] Wang & Gupta (2024), [4] Zeng et al. (2024),
[5] Gupta et al. (2025), [6] Gelfond & Lifschitz (1988), [7] Tudor et al. (2025)
The 16 desiderata for trustworthy AI are described in detail by Lenat & Marcus (2023), and so this section will only briefly summarize them. Table 1 indicates how various AI systems accomplish the desiderata for reference. For each of the 16+2 desiderata, we provide examples of how the s(CASP) system fulfills them. In some applications used for illustration, we describe how we successfully combine machine learning and s(CASP)-based automated commonsense reasoning with a goal of achieving human-level intelligence at least for domain-specific tasks.
@@ 1 Multiple Possible Worlds ! ASP Multiple Worlds[6]
#01 Explanation: Explainability is a commonly desired trait of AI, stating that any trustworthy AI must be able to explain any solution found. Its importance is reflected in the growing field of XAI.
@@
s(CASP): Goal-directed execution means that one poses queries that are answered by an SLD-resolution-like process, similar to Prolog, and therefore, a proof is explicitly constructed. This proof constitutes not only the explanation but also the justification of a given decision (positive or negative). In s(CASP), the justification for any query is the trace of the successful derivation, or in cases where the query fails, is the trace of the negated query. Moreover, the framework developed by Arias et al. (2020) allows the generation of justifications in natural language and navigable HTML files. Thus, explainability is native to s(CASP) and has been exploited in applications such as the physician advisor for chronic heart failure management by Chen et al. (2016) and by Forsante Oy (2025).
@@
Discussion: One of the great weaknesses of deep learning systems is that they are unexplainable black boxes, where it is difficult to justify a solution. The closest LLMs get is that they can be asked “why” an answer was given, but the explanation provided may be for a different alternative solution or the justification itself may be a hallucination.
@@
#02 Deduction: A foundational aspect of trustworthy AI is its capacity for human-like deduction, encompassing logical reasoning (e.g., modus ponens), arithmetic operations, exhaustive search, and the identification of contradictions or redundancies. This includes understanding logical connectives, in particular recognizing different forms of negation, i.e., not being able to conclude p is different from being able to conclude p is false.
s(CASP): Deduction is the basis of symbolic systems based on logic. Given premises p and p \$implies\$ q, we deduce q. Suppose we are given the premises that Tweety is a bird, bird(tweety), and the formula $\forall$$ X. bird(X) $\implies$$ flies(X). From these two premises, we can deduce that flies(tweety) holds, i.e., Tweety can fly. Constraint Programming adds declarative arithmetic processing to logic programming. The mortgage example by Holzbaur (1995) allows us to reason about the relation among the principal, P, the repayment rate, Mp, and the balance owing, B, i.e., the query ?- mortgage(P,12,,0.01,B,Mp) returns P = 6.14*R + 0.38*B. Finally, to deal with negation as failure (or default negation) when we lack evidence, the stable model semantics of Gelfond & Lifschitz (1988) is required, e.g., the rules p :- not q [^1] P :- Q expresses the premise Q \$implies\$ P. and q :- not p, have two stable models {p} and {q}. The stable model semantics can be realized via answer set programming (ASP). ASP also incorporates classical (or strong) negation in addition to default negation, e.g., -flies(X) :- penguin(X) means that X does not fly if it is a penguin (fly(X) is definitely false). Strong negation is used to give an explicit proof of a predicate’s falsehood.
@@
Discussion: Prior research has found that these deductive (along with inductive and abductive) reasoning skills are somewhat lacking in LLMs (Cheng et al., 2024; Abe et al., 2025). This is likely due in part to the fact that LLMs do not reason, but rather match patterns, and so are prone to confabulations. Cyc, on the other hand, provides mixed results. Cyc has over 1,000 reasoners for deductions that are each optimized for different kinds of knowledge, contexts, and solution finding. This provides a wide breadth of deduction ability, however there are disadvantages associated with having such a large number of reasoning engines. Namely, processing power required for running multiple reasoners, and overhead from meta-reasoners deciding which reasoners are needed. There is a trade-off between accuracy and speed associated with symbolic or neural methods.
#03 Induction: Effective inductive reasoning is essential for navigating complex, uncertain environments. It involves generalizing from specific observations, such as inferring species traits, and making temporal projections where the probability of a fact holding over time often follows predictable decay curves (e.g., linear, Gaussian). Such reasoning supports adaptive decision-making, despite the inherent risk of error. This type of reasoning is the foundation for explanation-based learning, which seeks to learn more with less data based on inducing general rules from specific examples.
s(CASP): Inductive Logic Programming (ILP), see survey by Cropper & Dumancic (2022), is a sub-field of machine learning that learns interpretable logic programs from labeled data and background knowledge. Unlike statistical models, ILP uses symbolic representations and logical inference to derive hypotheses that generalize observations in a human-understandable way. This makes ILP valuable in domains requiring transparency, prior knowledge integration, and formal reasoning, such as bio-informatics, legal systems, and explainable AI. The FOLD family of algorithms extends ILP by enabling the learning of default rules with exceptions, capturing common patterns of non-monotonic reasoning (Gupta et al. (2023); Wang & Gupta (2024)). Given the following data, the initially-learned rule is flies(X) :- bird(X) due to the data on “tweety”. Then from the example “pengu” FOLD learns that penguins are an exception to the rule, so the rule flies(X) is refined by adding not ab(X).
@@
bird(tweety). flies(tweety). bird(pengu). penguin(pengu). -flies(pengu).
@@
flies(X) :- bird(X), not ab(X). ab(X) :- penguin(X).
FOLD has been used to encode models learned from various real-world data into s(CASP) programs. All versions of the FOLD algorithm inductively learn rule-sets from datasets, provide a comparable accuracy to deep learning systems (particularly, the FOLD-SE system), and are completely explainable, thanks to the s(CASP)’s justification framework. The example above is a simple one for illustration purposes only, for more complete examples and formal review of using FOLD for ILP, please see Gupta et al. (2023) and Wang & Gupta (2024)
@@
#04 Analogy: Analogy is the ability to map similarity between objects of often disparate meaning. This is defined as a representational mapping from a known “source” domain to a novel “target” domain (Hall, 1989). In natural language, this commonly comes in the form of analogical phrases (“life is like a box of chocolates”). However, analogical reasoning is also used for various logic problems (such as Raven’s Progressive Matrices). Analogical reasoning has been historically difficult for AI systems.
@@
s(CASP): Symbolic systems have been used to solve the analogical reasoning problem with various degrees of success (Mitchell, 2021). Because these other successful systems (Falkenhainer et al., 1989; Hofstadter et al., 1995) are often predicated on turning analogical problems into predicates that can then be reasoned over, s(CASP) could be used as a more modern reasoner with at least as much success. In addition, analogies can be automatically generated by taking a partially grounded s(CASP) rule, then “lifting” it by replacing constants with variables, then instantiating these variables with different constants. Thus, a concrete procedure for boiling water can be used to generate an identical procedure for boiling milk. Work is in progress to automate such analogical reasoning using s(CASP).
@@
In the social chatbot application by Zeng et al. (2024), analogical reasoning has been realized by integrating s(CASP)’s reasoning with an LLM’s ability to convert text into knowledge represented as logic predicates. If the LLM-based text-to-predicate translation module recognizes that the current topic is about the movie Titanic, a s(CASP)-based analogical reasoner is triggered to find, for example, Catch Me If You Can movie as a candidate for the next discussion topic, since Leonardo DiCaprio leads in both.
relevant_topic(A, B, has_same_person’) :- movie(A), movie(B), person(C), act_in(C, A), act_in(C, B).
Additionally, LLMs can also help in generating analogical information, that is then used by the s(CASP) engine to gather analogous pieces of information:
@@
relevant_topic(’Batman Dark Knight Rises’, Titanic’, hero sacrifices himself for others’).
@@
Discussion: LLMs have been found to perform well at responding to natural language analogical phrases (Musker et al., 2024). However, in analogical reasoning, LLMs often make mistakes different than what humans would make due to a lack of deeper understanding. The ability to replicate responses to analogical queries is valuable but insufficient for true trustworthy AI. Alternatively, symbolic approaches can achieve good results on logical puzzles like RPMs, but lack the ability to integrate with natural language or make logical leaps like analogical phrases require without careful data population. In that area, Cyc can do well only on analogical phrases contained in its knowledge base.